Fermi National Laboratory
 

Operation Summary for October 24, 2011 - October 26, 2011

 

Monday October 24

The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with the MiniBooNE and NuMI experiments taking beam.

During the day shift, an RF person brought a MI RF station (MIRF17) back online at 8:43 AM. At 9:23 AM, MIRF17 tripped off; it reset. Operations established beam to the Meson FTBF experiment T-1008 at 10:04 AM. MI personnel began a g minus 2 (g-2) study at 1:38 PM.

During the evening shift, MI personnel completed their g-2 study at 4:52 PM. At 6:01 PM, BRF17 tripped off for the first of three times in 14 minutes; it reset. The NuMI beam permit tripped off at 7:20 PM on a RF incomplete gap indication, but there were no RF stations off; operators reset the beam permit, but this happened three more times; operator investigated. Operators reported at 8:42 PM that MIRF17 was causing the NuMI beam permit trips. Operators reported short down times due to problems with LRF5 and BRF14.

Tuesday October 25

The midnight shift began with Operations monitoring beam to the NuMI, MiniBooNE, and T-1008 experiments.

During the midnight shift, at 1:24 AM, the Booster West Anode Power Supply (WAPS) tripped off due to problems with BRF9; operators bypassed the station and reduced Booster intensity. Operations disabled beam to T-1008 at 2:05 AM. Operators reported at 2:24 AM that MIRF17 continued to cause NuMI beam permit trips; operators bypassed MIRF17. Operators reported at 3:12 AM that Booster was not responding well to the present RF configuration; operators loaded an energy conservation timeline into the Time Line Generator and called in a RF technician. At 4:09 AM, the Booster RF person reported that he was disconnecting BRF9 from the WAPS; a Booster person came in to help with tuning. At 7 AM, Neutron Therapy Facility (NTF) personnel began preparing to treat two patients: the first patient early on the day shift and the second one at noon. A RF technician brought MIRF17 back online at 7:53 AM.

During the day shift, at 8:33 AM, MIRF14 tripped off for the first of four times in 16 minutes; an RF tech investigated. Operators prepared the Antiproton Source (Pbar) rings for access at 8:48 AM to install equipment for future studies. Operations established beam to T-1008 at 10 AM. The Booster WAPS tripped off at 10:09 AM; an RF person reported that BRF10 had caused the trip. Booster RF personnel reported at 1 PM that they were bringing BRF9 online and connecting it to the WAPS. At 1:09 PM, operators reported that the NuMI beam permit dropped due to a temperature problem; NuMI personnel investigated. A NuMI person reset their humidifier and Operations reestablished beam to NuMI at 1:45 PM.

During the evening shift, at 6:28 PM, the NuMI beam permit tripped off on a RF fault and again there was no RF station off. Operators investigated and found that MIRF17 had gone to full voltage; this was the first of three similar trips in about two and half hours. Operators contacted a RF person and then turned off MIRF17 at 9:53 PM.

Wednesday October 26

The midnight shift began with Operations sending beam to the NuMI, MiniBooNE, and T-1008 experiments.

During the midnight shift, Operations disabled beam to T-1008 at 2:13 AM, this ended the experiment’s run at Fermilab for this year. Operations turned on the Pbar power supplies at 6:34 AM in preparation for a Mu2e study later in the morning.

Operations Coordinator (OpsCo) Plans for Today

The plans for today include Linac and Booster studies, Recycler proton studies, and Pbar reverse proton studies. A new Meson FTBF experiment, T-1015, will begin taking beam today.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/26/2011    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory